### v1

# End-to-End Visual Autonomous Parking via Control-Aided Attention

Abstract
Precise parking requires an end-to-end system where perception adaptively provides policy-relevant details—especially in critical areas where fine control decisions are essential. End-to-end learning offers a unified framework by directly mapping sensor inputs to control actions, but existing approaches lack effective synergy between perception and control. We find that transformer-based self-attention, when used alone, tends to produce unstable and temporally inconsistent spatial attention, which undermines the reliability of downstream policy decisions over time. Instead, we propose CAA-Policy, an end-to-end imitation learning system that allowes control signol to guide the learning of visual attention via a novel Control-Aided Attention (CAA) mechanism. For the first time, we train such an attention module in a self-supervised manner, using backpropagated gradients from the control outputs instead of from the training loss. This strategy encourages the attention to focus on visual features that induce high variance in action outputs, rather than merely minimizing the training loss—a shift we demonstrate leads to a more robust and generalizable policy. To further enhance stability, CAA-Policy integrates short-horizon waypoint prediction as an auxiliary task, and introduces a separately trained motion prediction module to robustly track the target spot over time. Extensive experiments in the CARLA simulator show that CAA-Policy consistently surpasses both the end-to-end learning baseline and the modular BEV segmentation + hybrid A* pipeline, achieving superior accuracy, robustness, and interpretability. Code is released at https://github.com/Joechencc/CAAPolicy.
I Introduction
Autonomous driving has traditionally been approached with modular pipelines that decompose the task into perception, prediction, and planning/control [1, 2, 3]. While such designs provide interpretability and modularity [4, 5, 6], they are prone to error accumulation across stages [7], redundant computation [6, 8], and reduced robustness under perturbations [9],
Fig. 1LABEL:fig:teasing(a). End-to-end (E2E) learning offers a unified alternative by mapping sensory inputs directly to control outputs through a single neural network [10, 11, 12], Fig. 1LABEL:fig:teasing(b).
Existing end-to-end (E2E) autonomous parking methods [13] typically predict control commands directly from multi-view images. While these approaches unify perception and control, their synergy remains weak: perception modules are trained solely through gradients backpropagated from task losses, and thus often fail to focus on control-relevant cues such as parking slot boundaries, lane markings, and obstacle gaps. This limitation is particularly pronounced in attention mechanisms. As shown in Fig. 2(a–b), a baseline without control guidance scatters attention across irrelevant regions and exhibits temporal inconsistency. Moreover, even when incorporating a automatic learned attention module such as CBAM[14], the model still fails to consistently emphasize near-target areas. Without stronger coupling between perception and control, attention tends to drift toward spurious features, resulting in misaligned perception and unreliable policy outputs.
To address these limitations, we propose CAA-Policy, an end-to-end autonomous driving framework that tightly couples perception and control (Fig. 1LABEL:fig:teasing(c)). At its core, it introduces a self-supervised Control-Aided Attention (CAA) mechanism that learns attention from control-output gradients rather than the loss gradients, enabling the attention module consistently focuses on control-sensitive regions near the target slot (Fig. 2(c)). Driving goals are further encoded as structured, interpretable tokens for goal-aware control, and short-horizon waypoint prediction is added as an auxiliary task to improve temporal consistency. Together, these components yield more accurate and robust autonomous parking.
Figure 2: Attention maps for target tracking at two consecutive frames ($t$ and t+1t{+}1): (a) without CAA, attention scatters over irrelevant areas; (b) with modified CBAM [14], an automatically learned attention module from task supervision (no auxiliary loss); (c) with CAA, attention focuses on the target and its vicinity; (d) ground-truth segmentation (yellow: target, blue: obstacles).
In summary, this paper makes the following contributions:
  • CAA-Policy, an end-to-end autonomous parking framework that tightly couples perception and control.
  • A novel Control-Aided Attention (CAA) mechanism that learn the attention model from control gradients wrt. perception features, making it consistantly focus on control-sensitive areas.
  • A learnable motion prediction module that tracks the vehicle’s current state using multiple historical motion frames for reliable current-step prediction.
  • A short-horizon waypoint prediction auxiliary task to enhance trajectory consistency and planning stability.
  • Extensive experiments demonstrate that CAA-Policy  significantly outperforms both modular hybrid A* and previous end-to-end baseline in trajectory accuracy, parking success rate, and interpretability in CARLA.
II Related Works
II-A Autonomous parking system
Prior work on autonomous parking ranges from traditional modular pipelines to modern End-to-End (E2E) methods. Early systems [15, 16, 17] typically combined occupancy mapping, hand-crafted planners (e.g., hybrid A* [18], RRT [19]), and rule-based controllers, offering interpretability but limited adaptability. Recent advances largely adopt either modular learning approaches or E2E learning paradigms [20, 13, 21].
Modular approaches follow the classic decomposition of perception, planning, and control, with each stage designed and tuned separately. Modular pipelines with hybrid A* planners [22, 23] face fundamental efficiency limitations that directly impact accuracy under practical trade-offs. Owing to the breadth-first-search nature of hybrid A*, planning becomes increasingly costly when the target is far, as the frontier space grows with each step. To mitigate the high computational burden, systems often reduce the frequency of re-planning with new perception inputs. However, this leads to blind execution based on outdated information, where occlusions near the target are ignored, frequently resulting in failure. This shortcoming has motivated the subsequent preference for learning-based policies and end-to-end approaches.
End-to-end approaches provide a unified alternative to modular pipelines by directly mapping sensory inputs to control outputs, making them particularly suitable for structured tasks like autonomous parking. Within this paradigm, one line of work formulates parking as a trajectory prediction task, where CNN–LSTM or Transformer architectures are supervised with trajectory waypoints rather than direct control commands [24, 25, 21]. Another line of work adopts imitation learning to predict low-level control actions directly. For example, E2EParking [13] maps multi-view images to control outputs, but encodes goal semantics only implicitly within segmentation labels, which limits generalization. To enhance goal awareness, later methods [21, 26] explicitly represent parking slots—either by introducing slot-map queries fused with BEV features or by applying cross-attention between goal points and perception features. Despite these variations, previous end-to-end methods lack explicit attention mechanisms to ensure perception focuses on critical regions for planning and control. We address this with Control-Aided Attention (CAA), which is trained to highlight control-sensitive areas.
II-B Deep Learning-based Motion Prediction
In our end-to-end parking setup, motion prediction is required to reliably track the ego vehicle’s state across multiple steps, ensuring that the planner operates on temporally consistent information. Prior end-to-end approaches such as E2EParking [13] employed segmentation-based tracking, but this design has been shown to fail in maintaining consistent predictions over multiple steps. Most traditional prediction methods are generally suitable for simple short-term tasks  [27, 28, 29], relying solely on the current state without modeling temporal dependencies. Recent temporal hybrid approaches combine multiple historical frames with physics-based constraints, employing architectures such as physics-informed RNN/LSTM [30, 31, 32], temporal CNN [33, 34], CNN–RNN hybrids [35, 36, 37], and attention mechanisms [38, 39, 40] to enhance long-horizon accuracy while maintaining physically feasible trajectories. Since our scenario is a static simulated environment, we adopt a lightweight CNN combined with a single-layer LSTM model. This design ensures model efficiency while leveraging temporal information to improve prediction generalization across different policies.
II-C Trajectory as Planning Output
Trajectory-based methods are common in autonomous parking, as they explicitly generate a feasible path from the vehicle’s pose to a target slot. Recent works integrate learning modules to enhance robustness and comfort, with most models [41, 25, 21] directly predicting trajectories from observations for controller execution. Alternatively, TCP [7] uses predicted trajectories as intermediate guidance for control. Building on this idea, CAA-Policy employs trajectory prediction as an auxiliary supervisory signal, benefiting from its structural guidance without requiring online replanning.
III Methodology
Figure 3: CAA-Policy consists of five main components: (1) Perception Backbone, (2) Feature Fusion, (3) Learnable motion prediction module, (4) Control-Aided Attention (CAA) module and (5) Prediction Heads. The prediction heads include control and waypoint prediction, while auxiliary heads for depth estimation and semantic segmentation are incorporated to improve feature learning. Notably, the Learnable motion prediction module is only used during inference, leveraging historical control and ego-motion information to reason about ego-vehicle dynamics and target position.
III-A Problem Statement
The goal of CAA-Policy is to control an autonomous vehicle to park safely and accurately into a designated slot using end-to-end learning. At each time step $t$, the vehicle receives the following inputs during inference:
  • Multi-view images $I_{t}\in\mathbb{R}^{6\times H\times W\times 3}$ captured from six surround cameras covering the vehicle: front, front-left, front-right, rear, rear-left, and rear-right views.
  • Current vehicle state $s_{t}$, including scalar speed $v_{t}$, planar acceleration $\mathbf{a}_{t}=[a_{x},a_{y}]$, and heading (yaw) $\psi_{t}$.
  • Target parking goal $g_{t}=(x_{t}^{\text{target}},y_{t}^{\text{target}},\psi_{t}^{\text{target}})$, specified in the global coordinate system and transformed into the vehicle’s ego frame at the start of the episode.
  • Historical motion information, comprising four consecutive frames, along with control commands $c_{t-4\dots t-1}$, ego-motion, and vehicle states $s_{t-4\dots t-1}$, is used by the motion prediction module.
The system outputs the future positions and control commands:
  • The vehicle’s next-frame position $(x_{t+1},y_{t+1})$, which is also used to update the ego-frame target coordinates for subsequent steps.
  • Short-horizon control commands $c_{t+1...t+4}=(\text{steering}_{t+1...t+4},\text{throttle}_{t+1...t+4},\text{brake}_{t+1...t+4})$ for the next 4 time step.
  • Short-horizon 2D waypoints for the next 4 time steps, used to ensure smooth and consistent future trajectory: $\mathbf{w}_{t+1:t+4}=[(x_{t+1},y_{t+1}),\dots,(x_{t+4},y_{t+4})]$.
  • Auxiliary predictions for perception tasks, including depth and BEV semantics.
The objective is to train a supervised policy $\pi$ from expert trajectories $\mathcal{D}=\{\tau^{n}\}_{n=1}^{N}$, where N os the number of expert trajectories, and each $\tau^{n}=\{(X^{n}_{t},Y^{n}_{t})\}_{t=1}^{T_{n}}$. Each trajectory $\tau^{n}$ consists of $T_{n}$ time steps with inputs $X^{n}_{t}$ and corresponding expert outputs $Y^{n}_{t}$ at each step:
  • $X^{n}_{t}$: all inputs (multi-view images, vehicle state, target goal, historical motion).
  • $Y^{n}_{t}$: expert labels for all supervised tasks (control, next pose, waypoints, depth, BEV semantics).
  • $T_{n}$: length (number of time steps) of trajectory $\tau^{n}$.
The policy $\pi$ is optimized to minimize
$\pi^{*}=\arg\min_{\pi}\frac{1}{N}\sum_{n=1}^{N}\frac{1}{T_{n}}\sum_{t=1}^{T_{n}}\mathcal{L}\left(Y_{t}^{n},\pi\left(X_{t}^{n}\right)\right),$
where $\mathcal{L}$ is a general supervised loss capturing all tasks.
Motion Prediction: Training vs. Inference. During training, the inputs do not require historical motion information, since ground-truth vehicle positions $\mathbf{w}_{t}$ and ego-target coordinates $\mathbf{g}_{0}$ are directly available and the motion prediction module is not used. In contrast, during inference, the module predicts these quantities from historical motion information to estimate the next-step vehicle state.
III-B Network Architecture
CAA-Policy consists of five main stages: (1) Perception Backbone, (2) Feature Fusion, (3) Learnable motion prediction module, (4) Control-Aided Attention(CAA) module, and (5) Prediction Heads, as illustrated in Fig. 3. Auxiliary heads for depth and semantic segmentation are also included. Notions are explicitly shown in Tab. I
TABLE I: Summary of major notations for CAA-Policy, grouped by module. The Symbol column has a light gray background for easy distinction, while module headers have a very light blue background.
Perception
$I_{t}$ Multi-view images at time $t$
$\mathbf{f}_{\text{BEV}}$ BEV feature from images
$\mathbf{f}_{\text{ego}}$ Encoded ego-motion
$\mathbf{f}_{\text{goal}}$ Encoded parking goal
$\mathbf{f}_{\text{fused}}$ Concatenated feature
Planning & Motion Prediction
$c_{t}$ Control command (steering, throttle, brake)
$\mathbf{w}_{t+1:t+4}$ Short-horizon future 2D waypoints
$(x_{t+1},y_{t+1})$ Predicted next position
$h_{t}$ Dynamics LSTM hidden state
$\bm{\mu}_{t},\bm{\sigma}^{2}_{t}$ Predicted 2D displacement mean & variance
Attention & Losses
AA Predicted BEV attention map (CAA)
$\mathbf{f}_{\text{CAA}}$ CAA-refined BEV feature
$\mathcal{L}_{\text{ctrl}},\mathcal{L}_{\text{wp}},\dots$ Task-specific losses
$\lambda_{\text{ctrl}},\lambda_{\text{wp}},\dots$ Loss weights
III-B1 Perception backbone
We adopt a ResNet-18 [42] encoder shared across six surround-view RGB images. The resulting features are fused into a unified bird’s-eye view (BEV) representation using the Lift-Splat-Shoot (LSS) framework [43]. To regularize the backbone and ensure the BEV features capture structural cues beyond control supervision, we employ auxiliary heads for semantic segmentation and depth estimation, ensuring the BEV features encode both geometric and semantic cues useful for downstream modules.
III-B2 Feature Fusion
In addition to the BEV feature extracted from multi-view images, ego-motion and parking goal information are required for downstream planning. Compared to previous end-to-end parking approaches [13], which embed the target slot implicitly into the BEV feature before tokenization, we explicitly tokenize all input information. The BEV features and ego-motion are encoded via CNN and MLP modules into compact representations, and the ego-target coordinates are explicitly tokenized. These representations are then concatenated to form the final fused feature:
$\mathbf{f}_{\text{BEV}},\mathbf{f}_{\text{ego}},\mathbf{f}_{\text{goal}}$ $=\mathrm{CNN}(\mathbf{I}_{1:N}),\mathrm{MLP}(\mathbf{e}),\mathrm{MLP}(\mathbf{g}),$ (1)
$\mathbf{f}_{\text{fused}}$ $=\mathrm{Concat}(\mathbf{f}_{\text{BEV}},\mathbf{f}_{\text{ego}},\mathbf{f}_{\text{goal}})$
where $\mathbf{I}_{1:N}$ denotes the set of $N$ multi-view images, $\mathbf{e}$ represents the ego-motion vector, $\mathbf{g}$ denotes the target parking slot coordinates in the ego-vehicle frame, $\mathbf{f}_{\text{BEV}}$, $\mathbf{f}_{\text{ego}}$, and $\mathbf{f}_{\text{goal}}$ are their corresponding tokenized features, and $\mathbf{f}_{\text{fused}}$ is the concatenated fused representation.
Target tokenization module(TTM): We explicitly convert the target parking slot coordinates into a token via $\mathbf{f}_{\text{goal}}=\text{MLP}(\mathbf{g})$. This explicit encoding preserves geometric information about the goal more effectively, improving trajectory prediction and control performance. The effect of this design is further analyzed in Sec. IV-D.
III-B3 Learnable Motion Prediction
We leverage the previous four frames, including control commands $c_{t-4:t-1}$ and vehicle states $s_{t-4:t-1}$, to estimate the current-step displacement relative to the last frame. Frame-wise features are encoded by a CNN and aggregated with an LSTM: $h_{t}=\mathrm{LSTM}\!\left(\{\mathrm{CNN}([c_{k},s_{k}])\}_{k=t-4}^{t-1}\right),$ where $h_{t}$ is the temporal latent state. Two MLP heads predict the mean and variance of the 2D displacement: $\bm{\mu}_{t}=W_{\mu}h_{t}+b_{\mu},\quad\log\bm{\sigma}^{2}_{t}=W_{\sigma}h_{t}+b_{\sigma}.$ The vehicle position is then updated as $(x_{t},y_{t})=(x_{t-1},y_{t-1})+\bm{\mu}_{t},$ which also refreshes the ego-frame target coordinates for planning.
During training, the predicted displacement distribution is supervised by the ground-truth displacement $\Delta\mathbf{x}_{t}^{\text{GT}}=(x_{t}^{\text{GT}}-x_{t-1}^{\text{GT}},\;y_{t}^{\text{GT}}-y_{t-1}^{\text{GT}})$. We minimize the negative log-likelihood (NLL) loss:
$\mathcal{L}_{\text{dyn}}=-\log\mathcal{N}\!\left(\Delta\mathbf{x}_{t}^{\text{GT}}\;\middle|\;\bm{\mu}_{t},\;\bm{\sigma}_{t}^{2}\right),$ (2)
where $\mathcal{N}(x\mid\mu,\sigma^{2})$ denotes the Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
Remark. This module is used only at inference to propagate states, while training directly uses ground-truth poses without optimizing this component.
III-B4 Control-Aided Attention (CAA)
The Control-Aided Attention (CAA) is introduced such that the attention module is learned to focus on control-sensive areas such that the perception module can provide more reliable information for the planning and control modules. CAA module predicts, for each BEV location, the expected control gradient in a self-supervised manner. Given fused perception features $\mathbf{f}_{\text{fused}}\in\mathbb{R}^{C_{b}\times X_{b}\times Y_{b}}$, it outputs a BEV-shaped attention map $\widehat{\nabla}_{f}\in\mathbb{R}^{X_{b}\times Y_{b}}$: $\widehat{\nabla}_{f}=\mathrm{MLP}(\mathbf{f}_{\text{fused}}),$ which refines the fused features via an element-wise product with channel-wise broadcasting:
$\mathbf{f}_{\text{CAA}}(c,i,j)=\mathbf{f}_{\text{fused}}(c,i,j)\cdot\widehat{\nabla}_{f}(i,j).$ (3)
III-B5 Control and Waypoint Prediction
Both control and waypoint prediction modules take the CAA-refined features $\mathbf{f}_{\text{CAA}}\in\mathbb{R}^{C_{b}\times X_{b}\times Y_{b}}$ as input. The control prediction head follows the structure of E2EParking [13], producing discrete control tokens autoregressively.
The waypoint prediction head explicitly models short-horizon 2D vehicle positions. Using only the current $\mathbf{f}_{\text{CAA}}$, it outputs future waypoints for the next few steps: $\mathbf{w}_{t+1:t+4}=\mathrm{WaypointDecoder}(\mathbf{f}_{\text{CAA}}),$ where $\mathbf{w}_{t+1:t+4}$ are the predicted waypoints in ego-centric coordinates.
This design allows both control and waypoint heads to leverage task-specific, CAA-refined perception features. The additional waypoint supervision provides explicit guidance on vehicle geometry and motion, improving trajectory stability and control accuracy.
III-C Loss Function
To achieve robust autonomous driving in complex urban scenes, CAA-Policy is trained under a unified multi-task framework that jointly optimizes planning, perception, and self-supervised attention alignment. Complementary objectives regularize each other: planning (control, waypoints) benefits from perception (segmentation, depth), while the Control-Aided Attention (CAA) loss aligns perception features with decision signals. The overall training objective integrates these components into a single loss:
$\mathcal{L}_{\text{total}}=\lambda_{\text{ctrl}}\mathcal{L}_{\text{ctrl}}+\lambda_{\text{wp}}\mathcal{L}_{\text{wp}}+\lambda_{\text{seg}}\mathcal{L}_{\text{seg}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{CAA}}\mathcal{L}_{\text{CAA}},$ (4)
where $\lambda$ is weighting coefficients balancing each task and each loss function is described as follows:
Control Loss
At each time step $t$, the ground-truth continuous control command $\tilde{c}_{t}^{\text{GT}}=(\text{steering}_{t},\text{throttle}_{t},\text{brake}_{t})$ is uniformly quantized into a discrete token index $c_{t}^{\text{GT}}$. Let $\hat{c}_{t}$ be the predicted token distribution over the discrete control vocabulary. The control loss is then defined as:
$\mathcal{L}_{\text{ctrl}}=\sum_{k=1}^{4}\text{CrossEntropy}(\hat{c}_{t+k},c_{t+k}^{\text{GT}}),$ (5)
where $c_{t+k}^{\text{GT}}$ denotes the ground-truth discrete control token at the $k$-th future step and $\hat{c}_{t+k}$ is the corresponding predicted probability distribution.
Waypoint Loss
At each time step $t$, the short-horizon future continuous waypoints $\tilde{\mathbf{w}}_{t+1:t+4}^{\text{GT}}=[(x_{t+1},y_{t+1}),\dots,(x_{t+4},y_{t+4})]$ are uniformly quantized into discrete token indices $\mathbf{w}_{t+1:t+4}^{\text{GT}}=[w_{t+1}^{\text{GT}},\dots,w_{t+4}^{\text{GT}}]$. Let $\hat{w}_{t+k}$ be the predicted token distribution for the $k$-th future waypoint. The waypoint loss is then defined as:
$\mathcal{L}_{\text{wp}}=\sum_{k=1}^{4}\text{CrossEntropy}(\hat{w}_{t+k},w_{t+k}^{\text{GT}}),$ (6)
where $w_{t+k}^{\text{GT}}$ denotes the ground-truth discrete waypoint token after quantization, and $\hat{w}_{t+k}$ is the predicted probability distribution.
Perception Losses
Following E2EParking [13], we supervise BEV semantic segmentation with cross-entropy $\mathcal{L}_{\text{seg}}=\text{CE}(\hat{M},M)$ and depth prediction with binary cross-entropy $\mathcal{L}_{\text{depth}}=\tfrac{1}{N_{\text{f}}}\sum_{i\in\mathcal{FG}}\text{BCE}(\hat{D}_{i},D_{i}),$ where $\hat{M}$ and $\hat{D}$ are predicted BEV semantic and depth maps, $M$ and $D$ are ground-truth maps, and $N_{\mathcal{FG}}$ is the number of foreground pixels $\mathcal{FG}$ with valid depth.
Control-Aided Attention (CAA) Loss
CAA module predicts a BEV-shaped attention map $A\in\mathbb{R}^{H\times W}$ that approximates the control sensitivity of the corresponding BEV location. Inspired by Grad-CAM [44], during training, the ground-truth control gradient is obtained via automatic differentiation: $\nabla_{f}^{\ast}=\partial\text{Control}/{\partial\mathbf{f}_{\text{fused}}}.$
The prediction is then supervised with an $L_{1}$ loss:
$\mathcal{L}_{\text{CAA}}=\left\|A-\nabla_{f}^{\ast}\right\|_{1}.$
This self-supervised loss encourages the perception module to focus on regions that most strongly affect downstream control predictions.
IV Experiments
Experiments are conducted in the CARLA 0.9.11 simulator (Town04-Opt), a parking lot scenario with 64 slots, following the extended environment of E2EParking [13]. Training and validation data are collected from successful parking trajectories generated by traditional modular pipelines with ground-truth BEV segmentation inputs, starting from random positions towards specified targets. For testing, we evaluate parking in unseen scenarios around novel parking slots. Model robustness is assessed using TSR, TFR, NTSR, NTFR, CR, and OR metrics, as reported in Tab. II.
TABLE II: Evaluation metrics for autonomous parking.
Metric Description
TSR Probability that the ego vehicle successfully parks into the target slot. Success: horizontal distance <0.6<0.6 m, longitudinal distance <1<1 m, orientation error $10^{\circ}$.
TFR Probability that the ego vehicle parks into the target slot but with unacceptable error.
NTR Probability that the ego vehicle parks into a non-target slot.
CR Probability of parking collision.
TR Probability that the ego vehicle fails to park within a specified time.
IV-A Implementation details
Most of our implementation follows E2EParking [13], including data collection, BEV feature size and the corresponding spatial resolution, backbone architecture, depth image parameters, encoder–decoder attention setup, and the optimization framework.
Compared to E2EParking, our setup mainly differs in dataset scale and training configuration. The training set contains $820$ trajectories (187,379187{,}379 frames) and the validation set includes $175$ tasks (40,22540{,}225 frames), both pre-collected for training the model. Each trajectory averages about $200$ frames, collected at $30$ Hz, while the model processes inputs at $0.1$ s intervals (every three steps). We train for $20$ epochs ($\approx$48 hours on 4 NVIDIA RTX 8000 GPUs with batch size $32$). Additionally, we use $6$ surrounding cameras (instead of $4$) to better cover blind spots. Testing is conducted in a separate environment by running complete parking episodes.
For expert demonstrations, rather than using human driving as in [13], we employ trajectories generated by a modular Hybrid A* planner, with control signals obtained from a bicycle model. Only successful and smooth trajectories are manually selected for training.
IV-B Baselines
We compare CAA-Policy against two baselines: E2E baseline [13] and Modular Hybrid A* baseline.
TABLE III: Comparison of parking performance across E2E baselines, ablations, the modular baseline, and the full model (Ours). All results are reported as $\mu\pm\sigma$ (mean $\pm$ standard deviation) in percentage over evaluation episodes. Metric definitions are provided in Table II.
[…]
Hybrid A* 59.1 $\pm$ 3.8 5.5 $\pm$ 2.2 0.0 $\pm$ 0.0 12.2 $\pm$ 2.9 21.1 $\pm$ 3.7
E2EParking (Reproduced) [13] 26.8 $\pm$ 5.1 3.5 $\pm$ 1.7 19.9 $\pm$ 3.4 34.2 $\pm$ 5.2 10.4 $\pm$ 3.2
E2EParking (Original) [13] 74.7 $\pm$ 5.0 1.0 $\pm$ 0.6 3.1 $\pm$ 1.3 12.0 $\pm$ 3.8 8.9 $\pm$ 3.3
CAA-Policy (Ablation) 18.7 $\pm$ 4.4 2.0 $\pm$ 1.1 20.3 $\pm$ 3.4 32.8 $\pm$ 4.8 18.0 $\pm$ 4.0
74.0 $\pm$ 5.7 1.8 $\pm$ 1.6 0.3 $\pm$ 0.2 13.5 $\pm$ 3.9 9.9 $\pm$ 3.2
76.0 $\pm$ 6.0 0.8 $\pm$ 0.5 0.0 $\pm$ 0.0 9.1 $\pm$ 3.6 12.2 $\pm$ 3.8
81.1 $\pm$ 4.5 4.5 $\pm$ 1.9 0.2 $\pm$ 0.2 4.4 $\pm$ 2.0 8.7 $\pm$ 3.2
80.2 $\pm$ 4.9 2.8 $\pm$ 1.6 5.9 $\pm$ 1.9 7.4 $\pm$ 2.8 2.1 $\pm$ 1.2
81.0 $\pm$ 5.4 2.3 $\pm$ 1.6 2.7 $\pm$ 0.8 6.2 $\pm$ 3.9 7.1 $\pm$ 3.4
CAA-Policy (Ours) 87.5 $\pm$ 3.7 2.3 $\pm$ 0.9 0.8 $\pm$ 0.6 3.5 $\pm$ 2.6 3.8 $\pm$ 2.2
IV-B1 E2E baseline study
We first examine the E2EParking [13] baseline, an end-to-end model predicting control commands from raw multi-view images. While the original reported TSR is 74.7%, it was obtained with a substantially larger training set and extended training schedule , which are not available for our experiments. Using only 820 expert trajectories and half the training time, our reproduction achieves a TSR of 26.8% with a Collision Rate of 34.2%, highlighting the sensitivity of end-to-end approaches to dataset scale: under limited data, E2EParking fails to generalize to unseen parking slots and cannot reliably track the target.
IV-B2 Modular baseline study
For comparison, we evaluate a modular autonomous driving pipeline using Hybrid A* [22, 23] for planning. Obstacles are derived from BEV segmentation, ego localization uses our motion prediction model, and a PID-controlled bicycle model handles vehicle control. This baseline achieves a TSR of 59.1% with zero NTSR, but exhibits a Timeout Rate of 21.1% and Collision Rate of 12.2%, indicating that modular pipelines, while stable, are limited in handling diverse or long-range parking scenarios.
IV-C Failure case study
Figure 4: Failure case study across different baselines and our method. Each row corresponds to a different experiment: (1) reproduced E2EParking trained with limited data, (2) Modular Hybrid A* baseline, (3) original E2EParking checkpoint trained with large-scale data and longer schedule, (4) CAA-Policy (Ours). Each column shows: (a) distribution of key metrics (TSR, TFR, NTSR, CR, TR) as pie charts, (b-d) the top three representative failure modes observed in that experiment. This layout allows direct comparison of overall performance and failure types across baselines and our method. Note that Trajectories are color-coded according to time sequence. For the Modular Hybrid A* baseline, red trajectories indicate the planned GT trajectory, which the controller struggles to follow on consecutive turns.
E2EParking: In Fig. 4, we analyze both the reproduced version trained with limited data (first row) and the original checkpoint trained on large-scale data (third row). Our failure case analysis shows three dominant patterns: (i) collisions with nearby vehicles, (ii) misalignment to unintended slots, and (iii) timeouts where the vehicle fails to reach the target slot within the time limit. These issues primarily arise from the reliance of E2EParking on segmentation-based estimation of the current target, which serves as a single point of failure. Small inaccuracies in target estimation can propagate through motion prediction, leading to larger trajectory deviations. In the low-data regime, the same failure types persist but occur more frequently, underscoring the sensitivity of purely end-to-end policies to training scale. These observations motivate our introduction of explicit modules (e.g., TTM and waypoints) to provide stronger target awareness and trajectory guidance.
Modular Hybrid A*: As shown in the second row of Fig. 4, classical planning pipelines using Hybrid A* generally produce accurate ego and target estimates. The main factors behind residual failures are: (1) Compounding perception errors: minor inaccuracies in object detection or free-space estimation can misguide trajectory planning; (2) Online planning limitations: computing feasible paths entirely online in complex environments can be computationally intensive, limiting the update frequency of perception inputs; (3) Controller-induced tracking errors: executing planned trajectories via a separate controller (e.g., PID) can introduce deviations that accumulate over longer horizons. This is particularly noticeable in scenarios with consecutive sharp turns, where even small path errors can cause the vehicle to deviate significantly from the planned trajectory. Although Hybrid A* demonstrates robust planning capabilities, suboptimal or infeasible paths in these situations are challenging for controllers to track accurately. This motivates the design of CAA-Policy, which directly generates control commands rather than relying on complex path planning. By doing so, it reduces accumulated tracking errors.
CAA-Policy: Finally, we analyze CAA-Policy (fourth row of Fig. 4). Overall failure rates are significantly lower than other methods, with all error types controlled below 4%, and the few remaining errors occurring near the target. This demonstrates that the learnable motion prediction module effectively follows the vehicle’s current state, and the Target Tokenization Module (TTM) accurately abstracts the parking goal, together ensuring precise target tracking. The Waypoint module further improves temporal consistency of predicted motions, while the Control-Aided Attention (CAA) mechanism emphasizes control-sensitive regions, enhancing attention to target cues and increasing parking success.
IV-D Ablation Study
We analyze the contribution of each key module in CAA-Policy - Control-Aided Attention (CAA), Target Tokenization Module (TTM), and Waypoint Prediction Head (Waypoints) - and their combinations to quantify complementary benefits and overall performance gains shown in Table III.
IV-D1 CAA module
The CAA module refines perception features to emphasize control-relevant regions. Applied alone, it performs worse than the baseline (18.7% TSR vs. 26.8%), as the backbone alone may provide insufficient guidance, causing reinforcement of incorrect attention patterns. This results in higher non-target success (20.3%) and a large collision rate (32.8%), highlighting that CAA alone cannot reliably guide the model without complementary trajectory information.
IV-D2 Target Tokenization Module (TTM)
The TTM module raises TSR from 26.8% (baseline E2EParking) to 74.0%. By explicitly encoding the target parking slot, it clarifies the intended goal for the agent, stabilizing control and reducing ambiguous behaviors. It also mitigates drifting into non-target slots, with NTSR dropping from 19.9% to 0.3%. Nevertheless, some failures remain, as reflected in the collision rate (13.5%) and timeout rate (9.9%), indicating that TTM further benefits from complementary modules providing trajectory guidance.
IV-D3 Waypoint module
Applying the Waypoint module alone, the model achieves a TSR of 76.0%, a substantial improvement over the reproduced baseline (26.8%) and comparable to the original E2EParking trained on large-scale data, while reducing the Collision Rate to 9.1%. This demonstrates that explicit intermediate waypoints provide strong spatial guidance, constraining the agent’s trajectory and reducing unintended deviations. However, the Timeout Rate remains relatively high (12.2%).
IV-D4 Module Synergy
We investigate the synergy between TTM, CAA, and Waypoints (Table III). Removing any single module while keeping the other two yields TSR values around 80–81%, indicating complementary contributions. However, all partial combinations fall short of the full CAA-Policy, which achieves 87.5% TSR — roughly 6–7% higher than any ablation — showing that jointly integrating TTM, CAA, and Waypoints maximizes performance and reduces collisions, timeouts, and non-target successes.
V Limitations
Despite CAA-Policy’s strong performance, limitations remain: evaluation excludes dynamic objects, limiting assessment in interactive scenarios; robustness to speed and target noise is partial, and environmental variations such as lighting and weather remain unaddressed; finally, as experiments are in simulation, real-world transfer requires further study.
VI Conclusion
In this work, we present CAA-Policy, an end-to-end autonomous parking framework that integrates Control-Aided Attention (CAA) for task-specific perception refinement, a Target Tokenization Module (TTM) for explicit target position representation, and a Waypoint Prediction auxiliary task to enforce temporal consistency. By explicitly predicting both vehicle motion and target location, and leveraging a short-horizon vehicle motion prediction model, CAA-Policy achieves robust and accurate tracking of ego and target positions in complex environments. Extensive experiments in CARLA demonstrate that each module contributes positively to performance, and their combination leads to improved trajectory stability, control accuracy, and overall success rates. These results highlight the effectiveness of combining perception refinement, temporal reasoning, and explicit motion supervision in end-to-end autonomous driving systems.
Appendix
In the supplementary material, we provide additional visualizations and analyses including: (1) a comparison of different target tracking strategies, (2) an evaluation of computation cost between end-to-end and modular baselines, and (3) a robustness study under perturbations in vehicle speed and target noise.
-A Target Tracking comparison
We evaluate approaches for tracking both ego and target positions: (1) a segmentation-based method [13], (2) a motion prediction model using a single historical frame, and (3) a motion prediction model leveraging multiple historical frames.
TABLE I: Target tracking comparison using Average Displacement Error (ADE) and variance. Inference time is negligible.
Method FDE (m) $\downarrow$ Variance $\downarrow$
Segmentation-based 0.082 0.127
Dynamics (single-frame) 0.403 0.020
Dynamics (multi-frame) 0.081 0.002
As shown in Tab. I, the segmentation-based method achieves moderate average error (FDE 0.082m) but is highly unstable, with variance reaching 0.127. This instability stems from its reliance on frame-by-frame segmentation: a single mis-segmented frame can lead to rapid divergence in subsequent motion and target predictions. Besides, the single-frame dynamics model cannot generalize well, resulting in a large FDE of 0.403m despite relatively low variance. In contrast, the multi-frame dynamics model achieves both the lowest error (0.081m) and minimal variance (0.002), as incorporating multiple historical frames smooths out noise and captures more reliable motion patterns, with negligible additional computational cost.
-B Computation cost: E2E vs Modular
TABLE II: Computation cost comparison between E2EParking and modular baselines. Inference time refers to the average time to execute a full trajectory. For E2EP methods, GPU, CPU, and runtime metrics are averaged across all end-to-end variants for clarity.
[…]
E2EParking (Ours) 2182 MB 6.1 GB 118.25
Modular: Hybrid A* 2092 MB 6.6 GB 188.75
Table II reports the average memory usage, CPU load, and inference time for E2EParking and the modular Hybrid A* method. All end-to-end variants exhibit similar GPU, CPU, and runtime profiles, which are summarized as an average for clarity. Compared to the modular approach, E2EParking uses slightly more GPU memory but substantially less CPU, completing inference faster due to its single-pass end-to-end design. These results highlight the practical efficiency of end-to-end methods for real-time deployment.
-C Resiliency to Speed and Target Noise
We evaluate the robustness of CAA-Policy against the modular Hybrid A* method under perturbations in vehicle speed and target position, with noise applied as a standard deviation around the ground truth values. As shown in Tab. III, end-to-end methods are generally more resilient to both types of noise, particularly for target position, likely due to explicit target augmentation during training. In contrast, learnable dynamics models are less robust to speed noise, resulting in larger performance drops. Overall, while all methods exhibit proportionally higher sensitivity with increasing noise levels, CAA-Policy consistently maintains superior TSR stability, highlighting the benefits of its end-to-end design and target-aware training.
TABLE III: Comparison of end-to-end and modular methods under vehicle speed and target position noise. TSR changes indicate the relative difference compared to the no-noise scenario.
[…]
Modular Hybrid A* 10 -0.8% 0.2 -0.3%
Modular Hybrid A* 30 -4.9% 0.5 -1.3%
Biguide 10 -0.2% 0.2 +0.2%
Biguide 30 -3.5% 0.5 -0.0%